Determine the refracted index of a certain piece of glass if the speed of light in it is 1.90x10^8 m/s

1
Expert's answer
2022-05-23T10:17:44-0400
"n=\frac{c}{v}"

"n=\frac{3.00*10^8}{1.90*10^8}=1.58"